# An Empirical Study of LLM-as-a-Judge for LLM Evaluation: Fine-tuned Judge Models are Task-specific Classifiers

### v1

@@ 3 Conclusion
In this work, we empirically study on the judge models for LLM evaluation. As revealed in our experiments, despite achieving superior evaluation performance on the in-domain testset, the fine-tuned judge models underperforms GPT4 in terms of generalibility and fairness by a large margin, which we believe cannot be amended by simply finetuning on more evaluation data. It is advisable to exercise caution when leveraging fine-tuned judge models for evaluation in real applications, watching for the overlap between the evaluation scenario and the training process.
@@
Our work still has some limitations: 1) Due to the lack of related work, we primarily relied on cross-validation to assess the generalizability of the four fine-tuned judge models. To conduct a more thorough evaluation of the generalizability, it would be beneficial to incorporate additional independent testsets encompassing a broader range of evaluation schemes, such as reference augmentation and domain-specific evaluation. 2) The work of Zeng et al. (2023) is only a general assessment of evaluator bias, and we did not include fine-grained assessment for different biases, such as formality bias, verbosity bias, etc. 3) Due to time and resource constraints, we did not incorporate manual inspection into the meta-evaluation process. Including human evaluators would enhance the credibility of our claims.
@@ A.1 Training Settings
As mentioned in Section 2.1, we fine-tune our own judge models based on the four groups of data (JudgeLM Zhu et al. (2023b), PandaLM Wang et al. (2024), Auto-J Li et al. (2023a), Prometheus Kim et al. (2023)), both in generation-style and in classification-style, for the purpose of comparison. We train all the models on NVIDIA A100-80GB GPUs with Huggingface-transformers Wolf et al. (2020) and DeepSpeed Rasley et al. (2020). Detailed hyper-parameters are presented in Table 7. Notice when comparing generation and classification models, we adopt the same prompt template and same hyper-parameters, with the only difference lies in the prediction method. For generation model, the prediction head reused the pretrained language model head, and is trained akin to the process of language modeling. For classification (regression) model, the prediction head is newly initialized as a linear projection layer, and is decoupled from the language modeling process [^5] Please refer to the class AutoModelForSequence Classification in Huggingface library for more details., as shown in Figure 4.
@@ A.2 Prompt Templates
As mentioned in Section 2.1, 2.2 and 2.3, we take the publicly released checkpoints of the four models and validate their performance. We use the same prompt templates as defined by the four open-source models, as presented from Figure 5 to Figure 12. For JudgeLM and PandaLM, their predefined prompts are in the form of pairwise selection, and we make slight modifications to apply them on pointwise grading. For Prometheus, the predefined prompt is in the form of pointwise grading, and we make slight modifications to apply it on pairwise selection. For Auto-J, they predefined prompts both for pairwise selection and pointwise grading.
Configuration Vicuna-based DeBERTa-based
base model Vicuna-7B DeBERTaV3-large
@@
learning rate schedule cosine decay cosine decay
@@
Table 7: Configurations of the judge models fine-tuned by us in Section 2.1. Both classification and generation judge models leverage the same group of configs based on their foundation model.

### v2

@@ 5 Conclusion
In this work, we conduct an empirical study on the judge models for LLM evaluation. As revealed in our experiments, despite achieving superior evaluation performance on the in-domain testset, the fine-tuned judge models underperforms GPT-4 in terms of several aspects by a large margin, which we believes originates from the task-specific fine-tuning process. We also propose an indicator to quantify the reliability of fine-tuned judges.
Although it is possible to incorporate more diverse fine-tuning data to amend the limitations, as the potential of LLM extends beyond boundaries, there would always be new domains and tasks not covered by the fine-tuning process. Therefore, to draw a conclusion, the fine-tuned judge model cannot serve as a general substitution for GPT-4 in terms of LLM evaluation. It is advisable to exercise caution when leveraging fine-tuned judge models for evaluation in real applications, watching for the overlap between the evaluation scenario and the fine-tuning process.
@@
Our work still has some limitations: 1) The reliability score proposed in our paper can be combined with human inspection or model ensemble, to further improve the LLM evaluation pipeline. For example, when applying judge model for LLM evaluation, we can select the least reliable samples to be evaluated by human evaluator, therefore improving the evaluation accuracy with minimal expense. Due to time and resource constraints , we have to leave this as future work. 2) The work of Zeng et al. (2023) is only a general assessment of evaluator bias, and we did not include fine-grained assessment for different biases, such as position bias Wang et al. (2023a), verbosity bias Saito et al. (2023), etc. 3) Due to time constraints, we did not incorporate manual inspection into the meta-evaluation process. Including human evaluators would enhance the credibility of our claims.
@@ A.1 Training Settings
As mentioned in Section 3, we fine-tune our own judge models based on the four groups of data (JudgeLM Zhu et al. (2023b), PandaLM Wang et al. (2024), Auto-J Li et al. (2023a), Prometheus Kim et al. (2023)), both in generation-style and in classification-style, for the purpose of comparison.
We train all the models on NVIDIA A100-80GB GPUs with Huggingface-transformers Wolf et al. (2020) and DeepSpeed Rasley et al. (2020). Detailed hyper-parameters are presented in Table 10. Notice when comparing generation and classification models, we adopt the same prompt template and same hyper-parameters, with the only difference lies in the prediction method, as illstrated in Figure 5. For generation model, the prediction head reused the pretrained language model head, and is trained akin to the process of language modeling. For classification (regression) model, the prediction head is newly initialized as a linear projection layer, and is decoupled from the language modeling process [^9] Please refer to the class AutoModelForSequence Classification in Huggingface library for more details..
@@ A.2 Prompt Templates
As mentioned in Section 3, we take the publicly released checkpoints of the four fine-tuned judge models and validate their performance. To make a fair comparison, we make minimal modifications to their pre-defined prompts, to adapt them to different scenarios. The specific prompts designed for different sections are listed as follows:
  1. 1.
    For Section 3.2, we adopt the prompts presented in Figure 6 to 13 for cross validation. Notice for JudgeLM and PandaLM, their predefined prompts are in the form of pairwise selection, and we make slight modifications to apply them on pointwise grading. For Prometheus, the predefined prompt is in the form of pointwise grading, and we make slight modifications to apply it on pairwise selection. For Auto-J, they predefined prompts both for pairwise selection and pointwise grading. We also adopt the prompts presented from Figure 14 to 17 on MT-Bench, which are all adapted to multi-turn evaluation.
  2. 2.
    For Section 3.3, we adopt the prompts presented in Figure 6, 8, 10 and 12, as LLMBar is pair-wise selection.
  3. 3.
    For Section 3.4, we adopt the prompts presented in Figure 18 to 21 for JudgeLM, PandaLM and Auto-J, respectively. For Prometheus, as its original prompt comprises of scoring rubrics, we simply define the corresponding rubrics for different benchmarks.
  4. 4.
    For Section 3.5, we adopt the prompts presented in Figure 22 and Figure 23 for chain-of-thought prompting.
Configuration Vicuna DeBERTa
@@
scheduler cosine decay cosine decay
@@
Table 10: Configurations of the fine-tuned judge models. Both classification and generation models leverage the same group of configs based on their foundation model.
@@
Figure 14: Prompt template for JudgeLM applied for multi-turn grading.
Figure 15: Prompt template for PandaLM applied for multi-turn grading.
Figure 16: Prompt template for Auto-J applied for multi-turn grading.
Figure 17: Prompt template for Prometheus applied for multi-turn grading.
Figure 18: Prompt template for JudgeLM applied on SALAD-Bench.
Figure 19: Prompt template for Auto-J applied on HaluEval.
Figure 20: Prompt template for Auto-J applied on ToxicChat.
Figure 21: Prompt template for Auto-J applied on SALAD-Bench.
Figure 22: Prompt template for JudgeLM applied with chain-of-thought prompting.
Figure 23: Prompt template for PandaLM applied with chain-of-thought prompting.